# Reasoning Abilities of Large Language Models: In-Depth Analysis on the Abstraction and Reasoning Corpus

### v1

@@ Abstract.
The existing methods for evaluating the inference abilities of Large Language Models (LLMs) have been results-centric, making it difficult to assess the inference process. We introduce a new approach using the Abstract and Reasoning Corpus (ARC) dataset to evaluate the inference and contextual understanding abilities of large language models in a process-centric manner. ARC demands rigorous logical structures for problem-solving, making it a benchmark that facilitates the comparison of model inference abilities with humans. Experimental results confirm that while large language models possess weak inference abilities, they still lag in terms of logical coherence, compositionality, and productivity. Our experiments highlight the reasoning capabilities of LLMs, proposing development paths for achieving human-level reasoning.
@@ 1. Introduction
Recent Large Language Models (LLMs) have demonstrated performance levels close to that of humans, but experimental results showed that they lacked planning ability through thought or reasoning (Bubeck et al., 2023). Consequently, a key question in recent language model research is: Can LLMs think? To address this question, new benchmarks for measuring reasoning abilities such as MathVista (Lu et al., 2024), Bonagard-Logo (Nie et al., 2020), and Raven (Zhang et al., 2019) have been proposed. Among them, the Abstract and Reasoning Corpus (ARC) (Chollet, 2019) emerged to be one of the representative benchmarks for assessing reasoning abilities. As shown in Fig. 1 below, each task in ARC consists of 2–5 example pairs and a problem input grid. The goal is to infer rules from given example pairs and apply them to the problem input grid. Input and output grid size can vary from a minimum of to a maximum of , with each grid having up to 10 different colors.
@@
ARC remains an unsolved challenge despite its seemingly simple content and evaluation methods. It demands a high level of abstraction and multiple reasoning steps, reasons why conventional deep learning techniques have not achieved success. The best-performing models to date have only achieved an accuracy of 30% (Lab42, 2024), while LLMs have shown an accuracy of around 10% (Mirchandani et al., 2023). Compared to the average human accuracy of 80% (Johnson et al., 2021), these results suggest significant differences in reasoning and abstraction capabilities between humans and LLMs. However, in-depth research into how LLMs reason and how their reasoning differs from humans is lacking. This has led to calls for a shift from a results-focused evaluation to a more nuanced analysis of the process (Chang et al., 2024; Huang and Chang, 2023; Xue et al., 2023), indicating a need for a new perspective that evaluates reasoning abilities based on the process rather than just the outcome.
@@
To overcome the limitations of result-oriented analysis in artificial intelligence, this study embraces the Language of Thought Hypothesis (LoTH) (Fodor, 1975), which is predominantly applied in the philosophy of mind to explain human reasoning. According to LoTH, effective reasoning encompasses three essential characteristics: logical coherence, the ability to follow basic logical principles; compositionality, the capability to construct complex ideas from simpler components; and productivity, the capacity to formulate an indefinite number of thoughts or solutions using a finite set of elements.
While there have been studies questioning LLM’s reasoning abilities (Bubeck et al., 2023; Valmeekam et al., 2024), a detailed analysis is lacking. Therefore, this study aims to utilize ARC to examine their reasoning skills from three perspectives of LoTH. To achieve this goal, we designed three separate experiments as follows:
Logical Coherence: Using prompting techniques, LLMs are set to solve ARC tasks. By analyzing the types of ARC tasks it can solve and the process of solving them, we aim to determine whether LLMs are capable of logical reasoning and whether its logic is consistent.
Compositionality: ARC tasks are known to be solvable through the application of step-by-step functions (Hodel, 2023a). With such functions provided, we aim to ascertain whether LLMs can identify the combinations of functions that are needed to solve a task. This process can be broken down into two parts: understanding how LLMs manipulate the problem input and determining whether they can achieve the desired results through multiple steps of manipulation.
Productivity: We tested if LLMs can create new input-output pairs for ARC tasks. We selected tasks with multiple inputs leading to the same output, devised prompts for inverse transformation, and assessed LLMs’ ability to generate varied inputs based on these prompts.
@@
As a result, we have confirmed that the current level of LLM possesses a basic understanding of images and is capable of simple types of compositional object manipulations. However, compared to human reasoning abilities, LLM lags in three areas: 1) It is weak in understanding aspects such as objectness in images. 2) Its logical reasoning abilities, especially in a step-by-step manner, are weak. 3) It struggles with understanding and generating unseen representations.
@@ 2. Preliminaries
This section aims to explain why we chose the LoTH perspective and ARC before starting a detailed evaluation of LLM’s reasoning capabilities. First, we will look at existing definitions of reasoning abilities and show why LoTH is useful in the perspective of measuring intelligence in Section 2.1. Then, by looking at ARC benchmarks and how they vary in Section 2.2, we highlight the need for using both the building blocks of meaning and the rules of combining them to tackle ARC tasks. This approach matches the LoTH perspective, showing that ARC is a good benchmark for studying LLM through the point of view of human reasoning.
2.1. Assessing Reasoning Ability of LLMs
@@
The divergent claims about the abilities of LLMs stem from result-centric measurement methods. Turing was the first figure to shift the approach to inference towards consequential direction (Turing, 1950). Subsequently, Wiener (Wiener, 1950), McCulloch and Pitts (McCulloch and Pitts, 1943), and Rosenblatt (Rosenblatt, 1958) shifted to studying methods for measuring performance rather than focusing on the process. Recently, Chollet attempted to quantify inference abilities from a consequential perspective (Chollet, 2019). However, these studies all focus on what reasoning can achieve using a result-oriented approach, without specifying the elements that constitute reasoning ability. West et al. (West et al., 2024) raised concerns about evaluating the reasoning ability of LLMs from a consequentialist perspective, as the generation capability of LLMs may not necessarily depend on comprehension abilities.
Therefore, a new perspective is needed to evaluate AI’s inference processes; Language of Thought Hypothesis (LoTH) enhances discussions by integrating reasoning components with quantitative metrics. LoTH posits that inference involves manipulating mental representations, a view that dominates the philosophy of mind due to its explanatory power over logical coherence, compositionality, and productivity observed in human cognition. These mental representations are believed to have a compositional syntax and combinatorial semantics. Our study compares with prior works and evaluates LLMs’ inference capabilities through the LoTH, marking progress by assessing aspects like logical coherence, compositionality, and productivity.
@@ 2.2. Advantages of using ARC as Reasoning Benchmark
In exploring benchmarks suitable for evaluating inference abilities through the lens of the Language of Thought Hypothesis (LoTH), the Abstraction and Reasoning Corpus (ARC) emerges as a compelling candidate. ARC tasks, necessitating combinatorial syntax and compositional semantics, align with LoTH’s emphasis on logical coherence, compositionality, and productivity, validating its role as an appropriate benchmark for reasoning abilities.
2.2.1. Comparison Between ARC and Its Variants
When comparing ARC with its variant benchmarks, it becomes evident that ARC poses problems that require a combination to solve. Despite its simple rules, ARC remains a benchmark with relatively low accuracy. LLM achieves an accuracy of 15% (Qiu et al., 2024), while non-LLM AI models reach a maximum of 31% (Lab42, 2024), which significantly differs from the human average accuracy of 80% (Johnson et al., 2021). The excessively high complexity of abstraction and reasoning is a major challenge of ARC. Consequently, various attempts to lower the difficulty of ARC by securing intermediate stages have naturally emerged. Those variants mainly fall into two categories: one focuses on reducing the complexity of abstraction by simplifying input images, while the other aims to reduce the complexity of reasoning dimensions by decreasing the number of steps needed to solve. Below are examples of variant benchmarks.
1D-ARC (Xu et al., 2023b) is a benchmark that reduces the dimensionality of ARC from 2D to 1D, aiming to simplify the complexity of the ARC. This transformation simultaneously maintains the core prior knowledge of ARC while reducing the dimensionality, thus lowering the complexity of the existing ARC task for enhanced research feasibility. Most importantly, the one-dimensionality of 1D-ARC effectively tackles the challenge of object cohesion. The unresolved object cohesion, regarded as one of the fundamental aspects of human cognition, poses a challenge in solving ARC tasks. Due to its successful handling of object cohesion, LLMs exhibited high accuracy in solving 1D-ARC tasks.
MC-LARC (Shin et al., 2023) adopts a multiple-choice format, featuring a rule sentence derived from ARC tasks alongside similar, but incorrect rule sentences. Solvers are tasked with selecting the accurate rule from five textual options, using the ARC examples as a reference. Effective resolution of MC-LARC requires solvers to semantically link these examples to the text options and distinguish the correct choice from the false ones. The pivotal shift from ARC to MC-LARC is transitioning from generative tasks to multiple-choice tasks.
Mini-ARC (Kim et al., 2022) is a benchmark dataset limiting the grid size to 5 5. It simplifies the ARC task by maintaining the core prior knowledge of ARC. It is eliminating the unnecessary process of varying input/output grid sizes. However, unlike the previously introduced 1D-ARC and MC-LARC, Mini-ARC remains similar to ARC in that it retains the characteristics of a 2D generative task. As a result, Mini-ARC remains challenging similar to ARC.
GPT-4 showed strong performance in MC-LARC, with approximately 75% (Shin et al., 2023), and in 1D-ARC, with about 90% (Xu et al., 2023b). However, it exhibited lower performance in Mini-ARC, around 15% (Qiu et al., 2024). As shown in Fig. 6, MC-LARC and 1D-ARC reduced the difficulty on a reasoning step level, whereas Mini-ARC differs by reducing image complexity as a benchmark. The significant decrease in difficulty of benchmarks that reduced step-by-step reasoning complexity supports the idea that ARC tasks requires a combination of sequential transformations. The main difference between MC-LARC, 1D-ARC, Mini-ARC, and ARC lies in whether specific functions must be combined to solve the task. 1D-ARC is designed around relatively straightforward transformations, thereby obviating the necessity for devising complex sequences of transformation combinations. Similarly, MC-LARC provides multiple choices, easily solvable by comparing options without the need for multiple stages of inference. However, Mini-ARC presents a unique challenge. Although the grid size is fixed at , the steps required to solve the task as complex as those in the standard ARC. Therefore, the performance difference between MC-LARC, 1D-ARC, and Mini-ARC implies the necessity of combinatorial syntax to combine functions to solve ARC.
2.2.2. Improvements on ARC Accuracy By Adding Object Information
ARC is a benchmark that requires extracting a large amount of semantic information from given images. Two prior research findings support this: 1) the improvement observed in studies providing additional information to solve ARC, and 2) comparative research results with other benchmarks. In one experiment, incorporating object information represented as graphs led to a notable increase in accuracy, with results nearly doublin (Xu et al., 2023b). This significant increase in accuracy underscores the importance of semantic information as a critical factor in resolving ARC tasks. Further research indicate that ARC contains more abstract information compared to other benchmarks, as illustrated in Table 1. Chollet, who proposed this benchmark, also argued that traditional feature extraction methods alone would not be sufficient to solve ARC, as it requires complex shapes and even additional information about transformation processes (Chollet, 2019). These observations confirm the importance of employing strategies that can interpret the abstract content essential to solving ARC challenges.
In summary, the variation in LLM accuracy with different ARC variations supports the need for combinatorial syntax in solving ARC, while the changes in accuracy with object information highlight the necessity of compositional semantics. These results demonstrate that ARC effectively represents the LoTH perspective, thereby validating the rationale for using it in this study.
@@ 3. Evaluating the Inferential Capabilities of LLMs Using the ARC Benchmark
To evaluate whether LLMs possess inferential capabilities, one could compare these capabilities to human reasoning. As explained in Section 2.1, according to the Language of Thought Hypothesis (LoTH), human reasoning can be broadly divided into three main components: Logical Coherence (Section 3.1), Compositionality (Section 3.2), and Productivity (Section 3.3). To examine each aspect of reasoning capabilities in LLMs, we utilized ARC, aiming to assess both the validity as a benchmark for evaluating reasoning capabilities and the inferential abilities of LLMs themselves.
@@ 3.1. Capability of LLMs 1: Logical Coherence
In Section 3.1, we aim to evaluate LLMs’ logical coherence, a first fundamental aspect of the LoTH. Logical coherence is the ability to understand a given logic and apply it consistently across different contexts. This concept is crucial in human cognitive processes as it facilitates the construction of sentence structures based on consistent logic, which is essential for solving various tasks. Such an ability is particularly relevant to the rule inference required in ARC tasks, where the challenge is to identify common logical patterns among given examples and use them to deduce the most logically coherent answer.
3.1.1. Experiment

### v2

@@ Abstract.
The existing methods for evaluating the inference abilities of Large Language Models (LLMs) have been results-centric, making it difficult to assess the inference process. We introduce a new approach using the Abstraction and Reasoning Corpus (ARC) dataset to evaluate the inference and contextual understanding abilities of large language models in a process-centric manner. ARC demands rigorous logical structures for problem-solving, making it a benchmark that facilitates the comparison of model inference abilities with humans. Experimental results confirm that while large language models possess weak inference abilities, they still lag in terms of logical coherence, compositionality, and productivity. Our experiments highlight the reasoning capabilities of LLMs, proposing development paths for achieving human-level reasoning.
@@ 1. Introduction
Recent Large Language Models (LLMs) have demonstrated performance levels close to that of humans, but experimental results showed that they lacked planning ability through thought or reasoning (Bubeck et al., 2023). Consequently, a key question in recent language model research is: Can LLMs think? To address this question, new benchmarks for measuring reasoning abilities such as MathVista (Lu et al., 2024), Bongard-Logo (Nie et al., 2020), and Raven (Zhang et al., 2019) have been proposed. Among them, the Abstraction and Reasoning Corpus (ARC) (Chollet, 2019) emerged to be one of the representative benchmarks for assessing reasoning abilities. As shown in Fig. 1 below, each task in ARC consists of 2–5 demonstration example pairs and a test example input grid. The goal is to infer rules from given example pairs and apply them to the test example. Input and output grid size can vary from a minimum of to a maximum of , with each grid having up to 10 different colors.
@@
The ARC remains an unsolved challenge despite its seemingly simple content and evaluation methods. It demands a high level of abstraction and multiple reasoning steps, reasons why conventional deep learning techniques have not achieved success. The best-performing models to date have only achieved an accuracy of 3-40% (Lab42, 2024), while LLMs (GPT-4, PaLM) have shown an accuracy of around 10-20% (Mirchandani et al., 2023). Compared to the average human accuracy of 80% (Johnson et al., 2021), these results suggest significant differences in reasoning and abstraction capabilities between humans and LLMs. However, in-depth research into how LLMs reason and how their reasoning differs from humans is lacking. This has led to calls for a shift from a results-focused evaluation to a more nuanced analysis of the process (Chang et al., 2024; Huang and Chang, 2023; Xue et al., 2023; Allen-Zhu and Li, 2023), indicating a need for a new perspective that evaluates reasoning abilities based on the process rather than just the outcome.
@@
To overcome the limitations of result-oriented analysis in artificial intelligence, this study adopts an existing theory on what constitutes a human’s reasoning ability. According to the Language of Thought Hypothesis (LoTH) (Fodor, 1975), human reasoning encompasses three essential characteristics: Logical Coherence, the ability to maintain consistency in reasoning; compositionality, the capability to construct complex ideas from simpler components; and productivity, the capacity to formulate an indefinite number of thoughts or solutions using a finite set of elements.
While attempts to evaluate logical coherence, compositionality, and productivity have existed before (Bubeck et al., 2023; Valmeekam et al., 2024), there were limitations in that the definitions of each component varied across papers and existing benchmarks showed insufficient performance in assessing each aspect. This study differs from previous research in two key ways: 1) by redefining concepts borrowed from psychology to fit the field of computer science, and 2) by evaluating all elements through the visual reasoning benchmark ARC. To achieve these goals, we have designed three separate experiments:
Logical Coherence: LoTH identifies two types of coherence. These are inferential coherence — the ability to apply logical reasoning across related instances coherently — and semantic coherence — the ability to maintain logical coherence in the reasoning process and results (Fodor and Pylyshyn, 1988). To verify both types of logical coherence, we augmented each solved ARC task with 100 similar test examples and evaluated the LLM’s performance on these related instances. Additionally, we analyzed the solution processes, identifying cases where correct answers were derived from flawed reasoning, to measure the LLM’s semantic coherence.
Compositionality: Compositionality refers to a system’s capacity to express one proposition is inherently linked to its ability to express related propositions (Fodor and Pylyshyn, 1988). In this study, we define compositionality as the ability to combine given semantics. Therefore, to evaluate compositionality, it is necessary to verify whether semantics can be combined as desired. Consequently, this study provided LLMs with step-by-step functions and examined whether they could identify the appropriate functions to solve ARC problems. Subsequently, we conducted an additional analysis to determine if the LLM could accurately predict the results from the given step-by-step functions and to understand the reasons for the failure.
Productivity: Productivity refers to the ability to infinitely create unseen expressions by combining a limited set of semantics (Fodor and Pylyshyn, 1988). However, it is difficult to quantitatively measure whether one can make an infinite number of unseen expressions. Therefore, previous studies have evaluated productivity by assessing whether rule-compliant unseen expressions can be created (van der Velde et al., 2004; Lake and Baroni, 2018; Hupkes et al., 2020). Similarly, in this study, to evaluate the ability to generate unseen expressions, we examined whether unseen ARC tasks that comply with the rules could be generated when given a set of functions.
@@
As a result, we have confirmed that the current level of LLM possesses a basic understanding of images and is capable of simple types of compositional object manipulations. However, compared to human reasoning abilities, LLMs lag in three areas: 1) It is not inferentially and semantically coherent. 2) Its logical reasoning abilities, especially in a step-by-step manner, are weak. 3) It struggles with understanding and generating unseen representations under complex constraints.
@@ 2. Preliminaries
This section aims to explain why we chose the LoTH perspective and ARC before starting a detailed evaluation of LLM’s reasoning capabilities. First, we will look at existing definitions of reasoning abilities and show why LoTH is useful in the perspective of measuring intelligence in Section 2.1. Then, in Section 2.2, we show that the ARC is an appropriate benchmark for studying LLMs from the perspective of human reasoning, as it 1) utilizes abstract semantics that can be generalized, and 2) is easy to modify.
2.1. Limitation on Assessing Reasoning Ability of LLMs
@@
The divergent claims about the reasoning abilities of LLMs stem from result-centric measurement methods. Turing was the first figure to shift the approach to inference towards consequential direction (Turing, 1950). Subsequently, Wiener (Wiener, 1950), McCulloch and Pitts (McCulloch and Pitts, 1943), and Rosenblatt (Rosenblatt, 1958) shifted to studying methods for measuring performance rather than focusing on the process. Recently, Chollet attempted to quantify inference abilities from a consequential perspective (Chollet, 2019). However, these studies all focus on what reasoning can achieve using a result-oriented approach, without specifying the elements that constitute reasoning ability. West et al. (West et al., 2024) raised concerns about evaluating the reasoning ability of LLMs from a consequentialist perspective, as the generation capability of LLMs may not necessarily depend on comprehension abilities.
Therefore, a new perspective is needed to evaluate AI’s inference processes; the Language of Thought Hypothesis (LoTH) enhances discussions by integrating reasoning components with quantitative metrics. LoTH posits that inference involves manipulating mental representations, a view that dominates the philosophy of mind due to its explanatory power over logical coherence, compositionality, and productivity observed in human cognition. These mental representations are believed to have a compositional syntax and combinatorial semantics. Our study compares with prior works and evaluates LLMs’ inference capabilities through the LoTH, marking progress by assessing aspects like logical coherence, compositionality, and productivity.
Logical coherence, compositionality, and productivity of LLMs have been evaluated separately in previous research. The ability of LLMs to perform deductive reasoning and maintain consistency in that reasoning has been considered an important evaluation metric. In this context, logical coherence is generally defined as the ability to construct coherent logic when solving specific problems (Zhang et al., 2024b). Meanwhile, whether LLMs possess compositional ability is one of the emerging questions. Given the lack of definition for compositionality, existing studies have not yet reached a unified opinion. Nevertheless, there have been attempts to define compositionality as the ability to understand and combine complex expressions, and to evaluate it by providing step-by-step prior knowledge necessary for solving problems requiring deductive reasoning (Lake and Baroni, 2018). Lastly, productivity is the ability to create new expressions from limited resources. With the emergence of generative AI, discussions are ongoing about what specific abilities or processes should be considered as productivity, and how to evaluate them. Some studies have judged this based on the accuracy and efficiency of the output generated using limited resources (van der Velde et al., 2004; Hupkes et al., 2020). However, these existing attempts have limitations in that the evaluation criteria and definitions vary across papers, and they have not compared the elements of reasoning to human reasoning processes.
@@ 2.2. Advantages of using ARC as Reasoning Benchmark
In exploring benchmarks suitable for evaluating inference abilities through the lens of the Language of Thought Hypothesis (LoTH), the Abstraction and Reasoning Corpus (ARC) emerges as a compelling candidate. ARC has two strengths compared to other reasoning benchmarks. First, it aligns with the LoTH perspective in that it can be solved through combinations of semantics. Second, it allows for easy task modification and generation, enabling flexible changes to objectives.
2.2.1. Core Properties of ARC
An important characteristic of the ARC benchmark is its requirement for extracting compositional semantics and combining them to resolve problems. This feature distinguishes ARC from other benchmarks and necessitates sophisticated problem-solving approaches. Two key research findings support this assertion:
Importance of Semantic Information: Studies demonstrating enhanced performance through supplementary information underscore the significance of semantic content in ARC task resolution. In a notable experiment, the integration of graph-represented object information resulted in a substantial accuracy increase, nearly doubling the success rate (Xu et al., 2023b). This marked improvement emphasizes the critical role of semantic information in effectively addressing ARC challenges.
High Abstraction Level of ARC: Comparative analyses reveal ARC’s higher abstraction level relative to other benchmarks, as evidenced in Table 1. Chollet, the benchmark’s proposer, contends that conventional feature extraction methodologies are insufficient for ARC, given its demand for complex shape interpretation and transformation process comprehension (Chollet, 2019). These findings highlight the necessity of employing advanced strategies capable of interpreting the abstract content integral to ARC problem-solving.
These observations confirm the importance of developing and implementing approaches that can effectively extract and utilize the complex, abstract information essential to solving ARC challenges. The unique nature of ARC necessitates a shift from traditional problem-solving paradigms toward more sophisticated, semantically-aware methodologies. Future research in this domain should focus on creating algorithms that can not only process visual information but also infer underlying rules, patterns, and transformations.
@@ 2.2.1. Core Properties of ARC
2.2.2. Flexibility in benchmark adaptation
Despite its simple rules, ARC remains a challenging benchmark with relatively low accuracy. LLMs achieve 15% accuracy (Qiu et al., 2024), traditional program synthesis models reach 26% (Wind, 2020), while human average accuracy is 80% (Johnson et al., 2021). To address this challenge, various ARC variants have emerged, focusing on reducing either abstraction complexity or reasoning dimensions.
Three notable ARC variants are:
1D-ARC (Xu et al., 2023b): Reduces dimensionality from 2D to 1D, simplifying complexity while retaining core knowledge. It effectively addresses object cohesion challenges, resulting in high LLM accuracy (about 90%).
MC-LARC (Shin et al., 2023): Adopts a multiple-choice format, transitioning from generative to selection tasks. GPT-4 showed strong performance (approximately 75%).
Mini-ARC (Kim et al., 2022): Limits grid size to 5x5, simplifying input while maintaining 2D generative characteristics. Performance remains challenging, similar to original ARC (around 15%).
These variations demonstrate ARC’s transformation flexibility and support the necessity of composition in solving ARC tasks. As shown in Fig. 3, MC-LARC and 1D-ARC reduced reasoning step complexity, while Mini-ARC focused on reducing image complexity. The performance differences among these variants imply that reducing the need for complex transformation combinations can significantly improve results, highlighting the importance of combinatorial syntax in solving ARC.
In summary, the ARC emerges as a compelling benchmark for evaluating inference abilities through the lens of the LoTH. ARC’s core strength lies in its requirement for extracting and combining compositional semantics to solve tasks, aligning well with the LoTH perspective. This is evidenced by improved performance when additional semantic information is provided, such as object information represented as graphs. The various ARC variants (1D-ARC, MC-LARC, and Mini-ARC) demonstrate their flexibility and adaptability for different experimental purposes. The significant performance differences among these variants, particularly the high accuracy in 1D-ARC and MC-LARC compared to the original ARC and Mini-ARC, support the necessity of combinatorial syntax and sequential transformations in solving ARC tasks. Furthermore, ARC’s high level of abstraction and reasoning complexity, as seen in its low accuracy rates for both LLMs and traditional models compared to human performance, underscores its value as a challenging benchmark. These characteristics collectively validate the rationale for using ARC in this study as an effective tool for exploring inference abilities in the context of the LoTH.
@@ 3. Evaluating the Inferential Capabilities of LLMs Using the ARC Benchmark
To evaluate whether LLMs possess inferential capabilities, one could compare these capabilities to human reasoning. As explained in Section 2.1, according to the Language of Thought Hypothesis (LoTH), human reasoning can be broadly divided into three main components: Logical Coherence (Section 3.1), Compositionality (Section 3.2), and Productivity (Section 3.3). We utilized Abstraction and Reasoning Corpus (ARC) to examine each aspect of the reasoning capabilities of LLMs from the perspective of LoTH.
@@ 3.1. Capability of LLMs 1: Logical Coherence
3.1.1. Motivation
Section 3.1 aims to evaluate the logical coherence of Large Language Models (LLMs). This is a fundamental aspect of the Language of Thought Hypothesis (LoTH), which considers coherence in two senses: inferential coherence and semantic coherence (Fodor and Pylyshyn, 1988). Semantic coherence refers to the ability to maintain logical consistency in the process and results of reasoning. Inferential coherence, on the other hand, is a system’s ability to consistently apply a specific type of logical inference across all relevant instances, given it can perform that inference in some cases. These concepts are crucial in human cognitive processes and relevant to the rule inference required in ARC tasks.
Our initial experiments primarily focused on measuring semantic coherence by evaluating whether the results produced by LLMs logically followed their problem-solving steps. This was done using various prompt techniques such as Chain of Thought (CoT) (Wei et al., 2022), Least to Most (LtM) (Zhou et al., 2023), and Tree of Thought (ToT) (Yao et al., 2023), similar to previous ARC solving attempts (Xu et al., 2023b; Mirchandani et al., 2023). We compared the coherence levels these different prompting strategies achieved, aiming to identify which techniques yielded the most semantically coherent results across diverse problem-solving scenarios. However, recognizing the limitations of this approach in addressing inferential coherence, we introduced supplementary experiments using augmented ARC tasks. These tasks, created through the Re-ARC program (Hodel, 2024), allow us to assess how consistently LLMs can apply logical patterns across variations of originally solved problems, providing a more comprehensive evaluation of their logical reasoning capabilities.
3.1.2. Comparison Across Prompting Techniques